### v1

# Assignment–Routing Optimization with Cutting-Plane Subtour Elimination: Solver and Benchmark Dataset

Abstract
We study a joint routing–assignment optimization problem in which a set of items must be paired one-to-one with a set of placeholders while simultaneously determining a Hamiltonian cycle that visits every node exactly once. Both the assignment and routing decisions are optimized jointly to minimize the total travel cost. In this work, we propose a method to solve this problem using an exact MIP formulation with Gurobi, including cutting-plane subtour elimination. With analysis of the computational complexity and through extensive experiments, we analyze the computational limitations of this approach as the problem size grows and reveal the challenges associated with the need for more efficient algorithms for larger instances. The dataset, formulations, and experimental results provided here can serve as benchmarks for future studies in this research area. GitHub repository : https://github.com/QL-YUAN/Joint-Assignment-Routing-Optimization
Keywords Joint Routing–Assignment Optimization $\cdot$ Mixed Integer Programming $\cdot$ Cutting-Plane Method $\cdot$ Datasets
1 Introduction
Many practical tasks in robotics, logistics, and manufacturing require transporting a collection of discrete objects from their current locations to designated target places. Such problems naturally couple three interdependent decision layers: (i) assignment — determining which object should go to which labeled placeholder; (ii) sequencing — deciding in what order the agent should pick up and deliver the objects; and (iii) routing or motion planning — finding a feasible path that minimizes travel cost while satisfying geometric and kinematic. This paper focus on first two concerns.
A range of realistic scenarios are directly related:
  • Vehicle-based pickup and drop-off. An autonomous vehicle or courier robot must pick up multiple passengers or packages (one at a time) and deliver each to a unique destination. Determining the optimal pickup–drop-off pairings and the overall route forms a cyclic tour that minimizes travel cost and maximizes efficiency.
  • Tabletop rearrangement with robot arms. A robotic manipulator moves scattered items (e.g., puzzle pieces or tools) into specific target positions on a workspace. Each item must be assigned to a placeholder, and the robot must plan sequence that visits all pick-and-place locations efficiently.
  • Automated packaging and assembly. A robotic or human operator sequentially transfers items (e.g., pastries, gadgets, or components) from a source area into predefined slots in packaging trays or boxes. The system must determine both which item goes where and in what order to perform the transfers to reduce time or effort.
  • Service robotics and delivery logistics. A service robot performs repeated pick-and-place tasks, such as collecting food orders or mail items and delivering each to a designated locker or customer. Optimizing the task sequence and assignments jointly improves response time and route efficiency.
Although compact in statement, this problem integrates the combinatorial structure of assignment and routing with precedence and connectivity constraints. The decision variables include a bijective mapping and a tour permutation, while the agent must obey unit-capacity and pickup-before-delivery constraints. The objective is typically to minimize total travel distance in this work, but extensions may incorporate other costs such manipulation costs, stopping overheads etc. which could be considered in futureworks.
1.1 Connections to existing problem classes
This problem lies at the intersection of several established families in operations research and robotics.
Pickup-and-delivery and the stacker-crane problem.
Pickup-and-delivery variants of the traveling salesman and vehicle routing problems (PDTSP, VRP-PD) enforce precedence constraints between pickup and delivery pairs under vehicle capacity limits. The stacker-crane abstraction (Hernández-Pérez and Salazar-González, 2004, 2007; Treleaven et al., 2012) is closely related, representing a single-vehicle system that must transport items between paired pickup and drop-off locations. These problems are NP-hard in general and have motivated a rich literature on branch-and-cut formulations, exact algorithms for small-scale instances, and heuristic or metaheuristic methods for larger ones.
Assignment–routing integration.
Classical formulations of the assignment problem (Kuhn, 1955) and the traveling salesman problem (Dantzig et al., 1954; Lawler et al., 1985; Applegate et al., 2006) address individual subcomponents of our task. However, when assignment and routing are coupled the search space expands combinatorially, requiring integrated formulations or hierarchical decomposition. Recent robotics literature has explored mixed-integer and sampling-based approaches for such coupled problems, yet scalability and global optimality remain open challenges. Modern solvers such as Gurobi (Gurobi Optimization, LLC, 2024) allow constraint families to be activated dynamically through cutting-plane method based callbacks, a strategy that has been successfully employed in subtour elimination for TSPs and connectivity-enforced routing (Applegate et al., 2006). This mechanism reduces model size and computational overhead by adding connectivity constraints only when violated, which we exploit in our implementation.
1.2 Purposes of the Article
We present a formulation that explicitly combines assignment and routing with precedence and connectivity constraints, forming a single Hamiltonian cycle over all pickup and delivery nodes.
The main purposes of this paper are:
  1. 1.
    A compact mixed-integer programming (MIP) model that jointly incorporates assignment, capacity, and cycle constraints, with an efficient Gurobi implementation using cutting-plane constraints to dynamically eliminate subtours.
  2. 2.
    Publicly available benchmark datasets, including input data and computed optimal solutions, to facilitate reproducibility and serve as a reference for future research.
  3. 3.
    Analysis of the computational complexity of this MIP formulation, highlighting the limitations of exact solution methods and motivating the development of more efficient heuristic algorithms.
2 Mathematical Formulation
Let $I=\{0,\dots,n-1\}$ be the set of items, $P=\{n,\dots,2n-1\}$ the set of placeholders, and $c_{ij}$ the cost of traversing edge $(i,j)$. Denote $V=I\cup P$ and $E=\{(i,j)\in V\times V:i\neq j\}$.
We define binary decision variables:
[…]
The objective is to minimize total traversal cost:
$\min\sum_{(i,j)\in E}c_{ij}x_{ij}.$
Constraints are as follows:
  1. 1.
    Degree: Each node has exactly two incident edges:
    $\sum_{j\in V,j\neq i}x_{ij}+\sum_{j\in V,j\neq i}x_{ji}=2,\quad\forall i\in V$
  2. 2.
    Assignment: Each item and placeholder is assigned twice:
    $\sum_{p\in P}a_{ip}=2,\quad\forall i\in I,\quad\sum_{i\in I}a_{ip}=2,\quad\forall p\in P$
  3. 3.
    Valid edges: Only item-to-placeholder edges are allowed, and edges can only exist if the assignment exists:
    $x_{ij}=0\text{ if both }i,j\in I\text{ or }i,j\in P,\quad x_{ip}\leq a_{ip},\quad x_{pi}\leq a_{ip},\forall i\in I,p\in P$
  4. 4.
    Fixed pair: In case of having starting and goal points, we specify starting position to be the last placeholder, goal point to be the last item, and enforce the last item to be assigned to the last placeholder as following constraints.
    $a_{n-1,2n-1}=1$
  5. 5.
    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method Kelley (1960); Westerlund and Pettersson (1995):
    $\sum_{i,j\in S,i\neq j}x_{ij}\leq|S|-1,\quad\forall\text{ subtours }S\subset V$
    This is enforced dynamically using subtour elimination callback function in Gurobi.
3 Computational Complexity
The pure assignment problem can be solved in polynomial time using the Hungarian algorithm Kuhn (1955) in $O(n^{3})$ complexity. However, the present formulation couples the assignment decision with routing optimization, the Traveling Salesman Problem (TSP) with fixed pick-place tasks constraints with complexity of $O(n!)$, making the problem significantly harder in $O(n!*n^{3})$ complexity.
3.1 Computational Complexity of The Proposed MIP Solver
The mixed-integer programming (MIP) solver in Gurobi employ a branch-and-bound framework augmented with cutting planes, preprocessing, heuristics and node presolve (commonly referred to as branch-and-cut). Specifically, the model requires finding both: (i) a bijective assignment between items and placeholders, and (ii) a single traversal order (tour) that visits all nodes exactly once while satisfying assignment consistency and pickup–delivery precedence constraints.
The MIP model contains the following variables and constraints:
  • $O(n^{2})$ binary routing variables $x_{ij}$,
  • $O(n^{2})$ binary assignment variables $a_{ip}$,
  • $O(n^{2})$ linear constraints.
  • An exponential number of subtour elimination constraints introduced dynamically by the cutting-plane procedure Basu et al. (2020).
Let $m$ denote the number of binary decision variables in the MIP. For our specific formulation, the dominant binary variables are the routing and assignment binaries:
$m=\Theta(n^{2}),$
At each search instance the solver typically (i) solves a linear programming (LP) relaxation and (ii) runs presolve / separation routines. Let $T_{\text{LP}}(s)$ denote the time to solve an LP of size $s$ (which is $O(2^{m})$) and $T_{\text{sep}}(s)$ denotes the time spent on all separation/cut generation at the instance. LP algorithms are proved to be polynomial in $s$ Karmarkar (1984), so we can write $T_{\text{LP}}(s)=\operatorname{poly}(s)$. However $T_{\text{sep}(s)}$ is exponential function of $s$ in worse-case, according to study in Basu et al. (2020). In other words, in worse-case, the MIP algorithm has exponential worst-case time complexity even for only one single instance. Thus, the worst-case computational effort grows exponentially with problem size ($n$), and only small to medium instances can be solved to proven optimality using current MIP solvers Basu et al. (2020); Optimization (2025).
4 Datasets
All the datasets files can be downloaded from project GitHub repository.
The following Table 1 shows a sample of data files in the datasets used for experiments. Here, Experiment is the experiment ID (starting from 1000), ID is the item/ placeholder index, pX, pY are the coordinates of the item, and tX, tY are the coordinates of the placeholders.
Table 1: Item and placeholder coordinates for Experiment 1000
Experiment ID pX pY tX tY
1000 0 0.521386 0.603842 0.974764 0.153932
1000 1 0.470942 0.203248 0.699089 0.447241
1000 2 0.528759 0.191036 0.017513 0.291025
[…]
Table 2: Number of node pairs and corresponding data files
Number of Node Pairs ($n$) File Name
10 experimental_n_10_data.csv
20 experimental_n_20_data.csv
32* experimental_n_32*_data.csv
100 experimental_n_100_data.csv
200 experimental_n_200_data.csv
300 experimental_n_300_data.csv
400 experimental_n_400_data.csv
500 experimental_n_500_data.csv
1000 experimental_n_1000_data.csv
Experiments and Results Files
All experiments were performed using the Gurobi Optimizer with a maximum runtime (timeout) set to sufficient time for completing each instance (When n=100 or smaller, 5mins is more than enough. However, when n=300, instances can easily take more than 10mins). Source code to run the experiment can be found in GitHub following the readme instruction. The solver was configured to utilize all available threads on the workstation, enabling parallel execution of the algorithm.
Tested dataset samples:
[…]
Table 4 provides a sample of the experimental results with each row being the result of one single experiment:
  • experiment_id: Unique identifier of the experiment.
  • best_cost: Best objective value found by Gurobi, which is the optimal solution cost.
  • dt: Total computation time in seconds. This should be within timeout settings.
  • assignment: Selected assignment solution (represented as a list of items assigned to placeholders).
  • edges: List of edges used in the solution (e.g., for routing problems).
Table 3: Tested problem instances, result files, and corresponding image folders
Instance Result File Image Folder
n=32_icn=32\_ic gurobi_n_32_ic_results.csv img_32_ic
n=32_xqn=32\_xq gurobi_n_32_xq_results.csv img_32_xq
$n=100$ gurobi_n_100_results.csv img_100
$n=200$ gurobi_n_200_results.csv img_200
$n=300$ gurobi_n_300_results_1_30.csv img_300
Table 4: Summary of Gurobi results for sample experiments (timeout = 5 min)
experiment_id best_cost dt [s] assignment edges
0 19.5360 9.532 […] […]
1 30.7337 10.894 […] […]
2 26.1106 2.873 […] […]
5 Experimental Results
Samples of the results are shown in Figures 1234, and summarized in Table 5. The complete dataset of experimental results is available in the accompanying GitHub repository for full reproducibility.
As evidenced in Table 5, the exact solver performs efficiently for small instances, consistently returning optimal solutions in under a second for problem sizes up to $n=32$. However, as the problem size increases — particularly for $n=300$ (30 sample instances are included in the dataset) — the computational time becomes significantly larger. For most instances at this scale, the solver requires more than 10 minutes to complete, even when parallelized.
These empirical findings highlight the inherent scalability limitations of exact Mixed-Integer Programming (MIP) formulations for combinatorial problems of this nature.
Such performance trends are consistent with the theoretical complexity of the problem: the formulation generalizes the Traveling Salesman Problem (TSP), which is well-known to be NP-hard. The worst-case computational complexity of branch-and-cut solvers applied to MIP formulations typically grows exponentially with the number of binary variables, which in this case is $m=\Theta(n^{2})$, due to pairwise interactions between nodes.
These results underscore a fundamental computational barrier: while exact methods are valuable for small instances or benchmarking purposes, their exponential growth in runtime renders them impractical for large-scale instances. Hence, developing and analyzing approximation algorithms or heuristics that deliver near-optimal solutions with significantly lower computational cost becomes not only practical but necessary. Such approaches can maintain high-quality solutions while scaling efficiently to larger values of $n$.
Table 5: Merged experimental results for $n=32$ XQ, $n=32$ Chess, $n=100$, and $n=200$ placeholders.
$n=32$ XQ $n=32$ Chess $n=100$ $n=200$
Item Best Cost Time (s) Item Best Cost Time (s) Exp ID Best Cost Time (s) Item Best Cost Time (s)
1 8.9811 0.3522 1 8.6045 0.2179 1 19.5360 9.6141 1 32.5257 25.6686
2 11.1315 0.0791 2 11.4452 0.0763 2 30.7337 10.9660 2 29.7790 206.0344
3 12.5431 0.2851 3 13.5874 0.2403 3 26.1106 3.7165 3 26.9493 47.0434
4 10.6235 0.0617 4 11.2942 0.2280 4 19.0954 24.1544 4 27.4590 49.4273
5 10.9402 0.2697 5 10.8882 0.3152 5 18.5475 7.0274 5 28.5458 124.0022
6 8.5998 0.3594 6 9.5426 2.2543 6 19.2728 14.9589 6 32.5525 78.4429
7 10.8217 0.1051 7 11.3253 0.2189 7 20.7782 20.2498 7 38.1046 95.0898
8 9.3054 0.6133 8 9.8737 0.1141 8 25.1340 2.5392 8 38.0598 45.8259
9 10.6951 0.3497 9 11.3153 0.4692 9 19.9261 6.4724 9 31.5440 65.1356
10 11.4083 0.0931 10 12.3702 0.1388 10 20.6597 10.5358 10 28.0293 106.9303
Average 0.3062 Average 0.3579 Average 11.5 Average 107.7101
(a) Experiment 1: Best path
(b) Experiment 2: Best path
(c) Experiment 3: Best path
(d) Experiment 4: Best path
Figure 1: First four sample results showing the computed optimal cycle for $n=32$ (xq).
(a) Experiment 1: Best path
(b) Experiment 2: Best path
(c) Experiment 3: Best path
(d) Experiment 4: Best path
Figure 2: First four sample results showing the computed optimal cycle for $n=32$ (ic).
(a) Experiment 1: Best path
(b) Experiment 2: Best path
(c) Experiment 3: Best path
(d) Experiment 4: Best path
Figure 3: First four sample results showing the computed optimal cycle for $n=100$.
(a) Experiment 1: Best path
(b) Experiment 2: Best path
(c) Experiment 3: Best path
(d) Experiment 4: Best path
Figure 4: First four sample results showing the computed optimal cycle for $n=200$.
6 Conclusion
In this work, we addressed a joint assignment–routing optimization problem, where item-to-placeholder pairings and a Hamiltonian tour are determined simultaneously to minimize total travel cost. We formulated the problem as a compact mixed-integer program (MIP) and solved it using Gurobi, employing a cutting-plane approach for subtour elimination to ensure tour feasibility.
Our experimental analysis confirms that the exact MIP solver performs well for moderate-sized instances (e.g., $n=32$), typically producing optimal solutions in fractions of a second. However, as the problem size increases, computational requirements grow rapidly: for $n=200$, solving times exceed 100 seconds; for $n=300$, runtimes often exceed 10 minutes per instance. These results highlight the scalability limitations of exact MIP-based methods in practice.
This study underscores the need for more scalable approaches — such as approximation algorithms, heuristics, or hybrid learning–search methods — to tackle larger problem sizes efficiently while maintaining high solution quality. Our publicly released dataset and benchmark results support reproducibility and offer a foundation for evaluating future algorithmic developments. By providing this benchmark, we aim to enable systematic comparison of solution quality, runtime, and empirical scaling across methods in future research.
Use of Generative AI.
This manuscript involved the use of generative AI tools (ChatGPT-4, OpenAI) during the drafting process. These tools were employed to assist with language refinement, equation formatting, and improving the clarity of technical explanations. All scientific content, methods, experimental results, and conclusions were developed and validated by the authors. Final manuscript versions were thoroughly reviewed to ensure accuracy and originality.

### v2

@@
 The dataset, formulations, and experimental results provided here can serve as benchmarks for future studies in this research area.
+Many pickup-placing with routing problems can be solved with such method with addition assignment constraints.
 GitHub repository : https://github.com/QL-YUAN/Joint-Assignment-Routing-Optimization
@@
 Pickup-and-delivery variants of the traveling salesman and vehicle routing problems (PDTSP, VRP-PD) enforce precedence constraints between pickup and delivery pairs under vehicle capacity limits.
-The stacker-crane abstraction (Hernández-Pérez and Salazar-González, 2004, 2007; Treleaven et al., 2012) is closely related, representing a single-vehicle system that must transport items between paired pickup and drop-off locations.
+The stacker-crane abstraction hernandez2004 ; hernandez2007 ; treleaven2012 is closely related, representing a single-vehicle system that must transport items between paired pickup and drop-off locations.
 These problems are NP-hard in general and have motivated a rich literature on branch-and-cut formulations, exact algorithms for small-scale instances, and heuristic or metaheuristic methods for larger ones.
 Assignment–routing integration.
-Classical formulations of the assignment problem (Kuhn, 1955) and the traveling salesman problem (Dantzig et al., 1954; Lawler et al., 1985; Applegate et al., 2006) address individual subcomponents of our task.
+Classical formulations of the assignment problem kuhn1955hungarian and the traveling salesman problem dantzig1954solution ; lawler1985traveling ; applegate2006traveling address individual subcomponents of our task.
 However, when assignment and routing are coupled the search space expands combinatorially, requiring integrated formulations or hierarchical decomposition.
 Recent robotics literature has explored mixed-integer and sampling-based approaches for such coupled problems, yet scalability and global optimality remain open challenges.
-Modern solvers such as Gurobi (Gurobi Optimization, LLC, 2024) allow constraint families to be activated dynamically through cutting-plane method based callbacks, a strategy that has been successfully employed in subtour elimination for TSPs and connectivity-enforced routing (Applegate et al., 2006).
+Modern solvers such as Gurobi gurobi2024manual allow constraint families to be activated dynamically through cutting-plane method based callbacks, a strategy that has been successfully employed in subtour elimination for TSPs and connectivity-enforced routing applegate2006traveling .
 This mechanism reduces model size and computational overhead by adding connectivity constraints only when violated, which we exploit in our implementation.
@@
 We define binary decision variables:
-[…]
+$x_{ij}=\begin{cases}1,&\text{if edge }(i,j)\text{ is included in the cycle},\\[4.0pt] 0,&\text{otherwise.}\end{cases}$
 The objective is to minimize total traversal cost:
 $\min\sum_{(i,j)\in E}c_{ij}x_{ij}.$
-Constraints are as follows:
+2.1 Simplified Model Constraints
+In order to solve the problem efficiently, in this problem modeling, only edges from item to placeholder is allowed. Constraints are as follows:
+  1. 1.
+    Valid edges: In order to make the model as simple as possible for high high computational efficiency , here only item-to-placeholder edges are allowed.
+Since straight forward post processing of edges can result in complete cycle with directions, such definition is fine for standard assignment and routing problem.
+The valid edges constraints are as follows:
+[…]
+  2. 2.
+    Assignment: Each item and placeholder is assigned twice:
+    $\sum_{p\in P}x_{ip}=2,\quad\forall i\in I,\quad\sum_{i\in I}x_{ip}=2,\quad\forall p\in P$
+  3. 3.
+    Fixed pair: In case of having starting and goal points, we specify starting position to be the last placeholder, goal point to be the last item, and enforce the last item to be assigned to the last placeholder as following constraints.
+    $x_{n-1,2n-1}=1,x_{2n-1,n-1}=0$
+  4. 4.
+    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method kelley1960cutting ; westerlund1995extended :
+    $\sum_{i,j\in S,i\neq j}x_{ij}\leq|S|-1,\quad\forall\text{ subtours }S\subset V$
+    This is enforced dynamically using subtour elimination callback function in Gurobi.
+2.2 Generalization of Modeling for Problem with More Constraints
+The formulation introduced above is suitable for the case of a single type of items and placeholders, where each item can be assigned to any placeholder.
+However, many practical applications involve multiple types or categories of items, each associated with a distinct subset of compatible placeholders.
+Examples include multi-type packing tasks, board or game piece arrangements, and robotic placement operations where items of the same type are interchangeable but must be placed in designated regions.
+We will have follow-up articals to discuss on methods to solve such problems.
+However, first of all, in such scenarios, the simplified model as mentioned above will have difficulty defining the constraints.
+In such cases we would like the edges to be directional instead of only connecting from item to placeholder.
+At the same time, we need to introduce additional binary decision variables:
+$a_{ip}=\begin{cases}1&\text{if item }i\text{ is assigned to placeholder }p\\ 0&\text{otherwise}\end{cases}.$
+Then, the constraints of the joint assignment-routing problem becomes:
   1. 1.
@@
   3. 3.
-    Valid edges: Only item-to-placeholder edges are allowed, and edges can only exist if the assignment exists:
+    Valid edges: Only item-to-placeholder edges are allowed, and edges can only exist if the assignment exists (here placeholder to item edges is allowed in this modeling):
     $x_{ij}=0\text{ if both }i,j\in I\text{ or }i,j\in P,\quad x_{ip}\leq a_{ip},\quad x_{pi}\leq a_{ip},\forall i\in I,p\in P$
@@
     $a_{n-1,2n-1}=1$
+    $x_{n-1,2n-1}=1,x_{2n-1,n-1}=0$
   5. 5.
-    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method Kelley (1960); Westerlund and Pettersson (1995):
+    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method:
     $\sum_{i,j\in S,i\neq j}x_{ij}\leq|S|-1,\quad\forall\text{ subtours }S\subset V$
-    This is enforced dynamically using subtour elimination callback function in Gurobi.
+  6. 6.
+    Following constraints need to be added to speed up the solver.
+    $x_{ij}+x_{ji}<=1,\forall i,\forall j,i\neq j$
 3 Computational Complexity
-The pure assignment problem can be solved in polynomial time using the Hungarian algorithm Kuhn (1955) in $O(n^{3})$ complexity.
+The pure assignment problem can be solved in polynomial time using the Hungarian algorithm kuhn1955hungarian in $O(n^{3})$ complexity.
 However, the present formulation couples the assignment decision with routing optimization, the Traveling Salesman Problem (TSP) with fixed pick-place tasks constraints with complexity of $O(n!)$, making the problem significantly harder in $O(n!*n^{3})$ complexity.
@@
 The MIP model contains the following variables and constraints:
-  • $O(n^{2})$ binary routing variables $x_{ij}$,
-  • $O(n^{2})$ binary assignment variables $a_{ip}$,
-  • $O(n^{2})$ linear constraints.
-  • An exponential number of subtour elimination constraints introduced dynamically by the cutting-plane procedure Basu et al. (2020).
-Let $m$ denote the number of binary decision variables in the MIP. For our specific formulation, the dominant binary variables are the routing and assignment binaries:
-$m=\Theta(n^{2}),$
+  • $O(n^{2})$ binary routing variables $x_{ij}$, and $O(n^{2})$ linear constraints.
+  • An exponential number of subtour elimination constraints introduced dynamically by the cutting-plane procedure basu2020complexitybranchandboundcuttingplanes .
+Let $m$ denote the number of binary decision variables in the MIP. For our specific formulation, the dominant binary variables are the routing and assignment binaries: $m=\Theta(n^{2}),$
 At each search instance the solver typically (i) solves a linear programming (LP) relaxation and (ii) runs presolve / separation routines.
 Let $T_{\text{LP}}(s)$ denote the time to solve an LP of size $s$ (which is $O(2^{m})$) and $T_{\text{sep}}(s)$ denotes the time spent on all separation/cut generation at the instance.
-LP algorithms are proved to be polynomial in $s$ Karmarkar (1984), so we can write $T_{\text{LP}}(s)=\operatorname{poly}(s)$.
-However $T_{\text{sep}(s)}$ is exponential function of $s$ in worse-case, according to study in Basu et al.
-(2020).
-In other words, in worse-case, the MIP algorithm has exponential worst-case time complexity even for only one single instance.
-Thus, the worst-case computational effort grows exponentially with problem size ($n$), and only small to medium instances can be solved to proven optimality using current MIP solvers Basu et al.
-(2020); Optimization (2025).
+LP algorithms are proved to be polynomial in $s$ karmarkar1984new , so we can write $T_{\text{LP}}(s)=\operatorname{poly}(s)$.
+However $T_{\text{sep}(s)}$ is exponential function of $s$ in worse-case, according to study in basu2020complexitybranchandboundcuttingplanes .
+In other words, in worse-case, the MIP algorithm has exponential worst-case time complexity.
+Thus, the worst-case computational effort grows exponentially with problem size ($n$), and only small to medium instances can be solved to proven optimality using current MIP solvers basu2020complexitybranchandboundcuttingplanes ; gurobi2025mipprimer .
 4 Datasets
-All the datasets files can be downloaded from project GitHub repository.
+All the datasets files can be downloaded from project GitHub repository. The experiments presented here applied the simplified model as introduced in Section 2.
 The following Table 1 shows a sample of data files in the datasets used for experiments.
@@
   • dt: Total computation time in seconds. This should be within timeout settings.
-  • assignment: Selected assignment solution (represented as a list of items assigned to placeholders).
   • edges: List of edges used in the solution (e.g., for routing problems).
@@
 Instance Result File Image Folder
-n=32_icn=32\_ic gurobi_n_32_ic_results.csv img_32_ic
-n=32_xqn=32\_xq gurobi_n_32_xq_results.csv img_32_xq
-$n=100$ gurobi_n_100_results.csv img_100
-$n=200$ gurobi_n_200_results.csv img_200
-$n=300$ gurobi_n_300_results_1_30.csv img_300
-Table 4: Summary of Gurobi results for sample experiments (timeout = 5 min)
+n=32_icn=32\_ic updated_experiment_n_32_ic_results.csv img_update_32_ic
+n=32_xqn=32\_xq updated_experiment_n_32_xq_results.csv img_update_32_xq
+$n=100$ updated_experiment_n_100_results.csv img_update_100
+$n=200$ updated_experiment_n_200_results.csv img_update_200
+$n=300$ updated_experiment_n_300_results_1_30.csv img_update_updated_300
+Table 4: Summary of Gurobi results for sample experiments (n=100)
 experiment_id best_cost dt [s] assignment edges
-0 19.5360 9.532 […] […]
-1 30.7337 10.894 […] […]
-2 26.1106 2.873 […] […]
+0 19.5360 […] […] […]
+1 30.7337 […] […] […]
+2 26.1106 […] […] […]
 5 Experimental Results
-Samples of the results are shown in Figures 1234, and summarized in Table 5. The complete dataset of experimental results is available in the accompanying GitHub repository for full reproducibility.
-As evidenced in Table 5, the exact solver performs efficiently for small instances, consistently returning optimal solutions in under a second for problem sizes up to $n=32$.
-However, as the problem size increases — particularly for $n=300$ (30 sample instances are included in the dataset) — the computational time becomes significantly larger.
-For most instances at this scale, the solver requires more than 10 minutes to complete, even when parallelized.
+Samples of the results are shown in Figures 1234, and summarized in Table 5.
+The complete dataset of experimental results is available in the accompanying GitHub repository for full reproducibility.
+As evidenced in Table 5, the exact solver performs efficiently for small instances, consistently returning optimal solutions in under a second for problem sizes up to $n=100$.
+However, as the problem size increases — particularly for $n=300$, the computational time reach one minutes.
+When n reach 1000, the solver often fails to provide a valid solution.
 These empirical findings highlight the inherent scalability limitations of exact Mixed-Integer Programming (MIP) formulations for combinatorial problems of this nature.
@@
 Such approaches can maintain high-quality solutions while scaling efficiently to larger values of $n$.
-Table 5: Merged experimental results for $n=32$ XQ, $n=32$ Chess, $n=100$, and $n=200$ placeholders.
-$n=32$ XQ $n=32$ Chess $n=100$ $n=200$
-Item Best Cost Time (s) Item Best Cost Time (s) Exp ID Best Cost Time (s) Item Best Cost Time (s)
-1 8.9811 0.3522 1 8.6045 0.2179 1 19.5360 9.6141 1 32.5257 25.6686
-2 11.1315 0.0791 2 11.4452 0.0763 2 30.7337 10.9660 2 29.7790 206.0344
-3 12.5431 0.2851 3 13.5874 0.2403 3 26.1106 3.7165 3 26.9493 47.0434
-4 10.6235 0.0617 4 11.2942 0.2280 4 19.0954 24.1544 4 27.4590 49.4273
-5 10.9402 0.2697 5 10.8882 0.3152 5 18.5475 7.0274 5 28.5458 124.0022
-6 8.5998 0.3594 6 9.5426 2.2543 6 19.2728 14.9589 6 32.5525 78.4429
-7 10.8217 0.1051 7 11.3253 0.2189 7 20.7782 20.2498 7 38.1046 95.0898
-8 9.3054 0.6133 8 9.8737 0.1141 8 25.1340 2.5392 8 38.0598 45.8259
-9 10.6951 0.3497 9 11.3153 0.4692 9 19.9261 6.4724 9 31.5440 65.1356
-10 11.4083 0.0931 10 12.3702 0.1388 10 20.6597 10.5358 10 28.0293 106.9303
-Average 0.3062 Average 0.3579 Average 11.5 Average 107.7101
+Table 5: Merged experimental results for $n=32$ XQ, $n=32$ Chess, $n=100$, $n=200$, and $n=300$ placeholders.
+$n=32$ XQ $n=32$ Chess $n=100$ $n=200$ $n=300$
+Item Best Cost Time (s) Best Cost Time (s) Best Cost Time (s) Best Cost Time (s) Best Cost Time (s)
+1 8.9811 0.1749 8.6045 0.0640 19.5360 0.5980 32.5257 1.8281 42.0017 76.0503
+2 11.1315 0.0184 11.4452 0.0204 30.7337 0.7149 29.7790 8.3149 33.1289 25.6637
+3 12.5431 0.2018 13.5874 0.0154 26.1106 0.8715 26.9493 4.2116 34.3303 44.0879
+4 10.6235 0.0195 11.2942 0.0401 19.0954 0.9859 27.4590 6.2985 35.1968 28.5358
+5 10.9402 0.0450 10.8882 0.0494 18.5475 2.0169 28.5458 14.5061 33.8118 77.8469
+6 8.5998 0.0611 9.5426 0.0869 19.2728 1.5554 32.5525 6.9940 37.4785 84.7612
+7 10.8217 0.0149 11.3253 0.0717 20.7782 0.8501 38.1054 8.7085 33.9530 26.5168
+8 9.3054 0.1165 9.8737 0.0133 25.1340 0.2294 38.0598 4.1589 40.6682 28.3229
+9 10.6951 0.0742 11.3153 0.0821 19.9261 0.2498 31.5435 10.0047 36.6766 51.3296
+10 11.4083 0.0331 12.3702 0.0123 20.6597 2.6301 28.0285 11.0655 38.6581 83.7323
+Average 0.0758 0.0452 0.9709 7.6993 52.2849
 (a) Experiment 1: Best path
@@
 We formulated the problem as a compact mixed-integer program (MIP) and solved it using Gurobi, employing a cutting-plane approach for subtour elimination to ensure tour feasibility.
-Our experimental analysis confirms that the exact MIP solver performs well for moderate-sized instances (e.g., $n=32$), typically producing optimal solutions in fractions of a second.
-However, as the problem size increases, computational requirements grow rapidly: for $n=200$, solving times exceed 100 seconds; for $n=300$, runtimes often exceed 10 minutes per instance.
+Many pickup-placing with routing problems, for example, arranging chess game (with different type of pieces) and packing food/ toys with different item patterns, can be solved with such method with addition assignment constraints without increasing the complexity of the problem.
+Our experimental analysis confirms that the exact MIP solver performs well for moderate-sized instances (e.g., $n=30$), typically producing optimal solutions in fractions of a second.
+However, as the problem size increases, computational requirements grow rapidly: for $n=300$, solving times reach 1 minute; for $n=1000$, solver often fails to work.
 These results highlight the scalability limitations of exact MIP-based methods in practice.